# Breaking the Gaussian Barrier: Residual-PAC Privacy for Automatic Privatization

### v2

@@ 4.1 Foundation of Residual-PAC Privacy
Consider any $f$-divergence $\mathtt{D}_{f}$, Theorem 1 of [46] shows that the posterior advantage $\Delta_{f}^{\delta}$ is bounded by the minimum $f$-divergence between the joint distribution of $(X,\mathcal{M}(X))$, denoted by $P_{X,\mathcal{M}(X)}$, and the product of the marginal distribution $P_{X}$ and any auxiliary output distribution $P_{W}$ independent of $X$:
@@
Corollary 2 follows from Theorem 1 of [46] and establishes that when $\mathcal{H}(X)$ is finite, residual privacy $\mathtt{R}_{f}^{\delta}$ is lower bounded by $\mathcal{H}(X|\mathcal{M}(X))-\mathtt{V}$, where $\mathtt{V}$ is independent of both data distribution $\mathcal{D}$ and mechanism $\mathcal{M}$. Since $\mathtt{V}$ is constant, $\mathtt{R}_{f}^{\delta}-\mathtt{V}$ effectively provides a privacy quantification lower-bounded by conditional entropy $\mathcal{H}(X|\mathcal{M}(X))$. More generally, inequality (11) holds without requiring $\mathcal{H}(X)<\infty$, provided $\mathtt{D}_{f}(\mathcal{D}\|\mathcal{R})<\infty$.
@@ 5.2 Directional-Selectivity of SR-PAC
Let $Z\in\mathbb{R}^{d}$ be an output vector produced by a deterministic mechanism $\mathcal{M}(X)$; throughout we assume $\Sigma_{Z}\succ 0$ and finite differential entropy $\mathcal{H}(Z)$. For any application, let $S_{\mathrm{task}}\subseteq\mathbb{R}^{d}$ denote a practitioner-chosen task-critical sub-space (the directions whose preservation matters most) and write $\Pi_{\mathrm{task}}$ for the orthogonal projector onto it.
Classification tasks. In what follows we illustrate the theory with multi-class classification, where $Z$ is the logit vector, $\hat{y}=\arg\max_{i}Z_{i}$, and $S_{\mathtt{lab}}:=\operatorname{span}\{e_{\ell}-e_{j}:\,j\neq\ell\}$, where $\mathtt{lab}$ means "label". Let $\Pi_{\mathtt{lab}}$ be the projector onto $S_{\mathtt{lab}}$. The analysis for a general $S_{\mathrm{task}}$ is identical after replacing $\mathtt{lab}$ by $\mathrm{task}$.
@@
For every unit vector $w$, let $g(w)\equiv\frac{1}{2}\mathtt{mmse}(\langle Z,w\rangle)$, where $\mathtt{mmse}(\langle Z,w\rangle)\equiv\mathbb{E}\Bigl{[}\bigl{\langle}Z,w\bigr{\rangle}\;-\;\mathbb{E}\bigl{[}\langle Z,w\rangle|Y\bigr{]}\Bigr{]}^{2}$ is the minimum mean-squared error of estimating the scalar random variable $\langle Z,w\rangle$ from the noisy observation $Y=Z+B$.
@@ 6 Experiments
We conduct two sets of experiments to evaluate our approach. First, we compare SR-PAC against Auto-PAC and Efficient-PAC (Appendix A) using CIFAR-10 [29], CIFAR-100 [29], MNIST [31], and AG-News [48] datasets, with results presented in Section 6.1. We use (R-)PAC to refer to the family of SR-PAC, Auto-PAC, and Efficient-PAC. Second, we extend this comparison to include DP by equalizing optimal posterior success rates of membership inference (Appendix C.2) across all four methods (SR-PAC, Auto-PAC, Efficient-PAC, and DP), making their privacy budgets comparable. For this comparison, we use Iris [17] and Rice [9] datasets, with results shown in Section 6.2. All experiments focus on output perturbation. Appendix S provides more details of the experiments.
@@
CIFAR-100 and Base Classifier. CIFAR-100 contains 50,000 training and 10,000 testing color images (each $32\times 32\times 3$), equally divided among 100 fine-grained classes (500 training and 100 testing images per class). Each image is converted to a $3\times 32\times 32$ tensor and normalized per channel to mean 0.5 and standard deviation 0.5 before being fed into the network. The unperturbed classifier is a deep convolutional neural network with three convolutional “blocks.” Each block consists of two $3\times 3$ convolutions (with BatchNorm and ReLU after each), followed by a $2\times 2$ max-pool, which sequentially maps inputs from $32\times 32$$16\times 16$$8\times 8$$4\times 4$, with channel widths increasing from 3 → 64 → 128 → 256. After flattening the resulting $256\times 4\times 4$ feature map into a 4096-dimensional vector, a three-layer MLP head (4096→512→256→100) with ReLU activations and 0.5 dropout between the first two fully connected layers produces a 100-dimensional logit vector. During training, this network minimizes cross-entropy loss over the CIFAR-100 classes; at inference, each normalized image is mapped to its 100-dimensional logits, and the predicted label is given by the $\arg\max$ of those logits. The unperturbed classifier achieves $0.5914\pm 0.0090$ accuracy.
@@
AG-News dataset and Base Classifier. AG-News comprises 120,000 training and 7,600 test articles equally divided among four classes (World, Sports, Business, Sci/Tech), i.e., 30,000 training and 1,900 test examples per class. Each example’s title and description are concatenated into one text string, then lowercased and split on whitespace (truncated or padded to 64 tokens). We build a 30,000-word vocabulary from the training split and map each token to its index (with out-of-vocabulary tokens as 0). Those indices feed into an nn.EmbeddingBag layer (embedding size 300, mean-pooling mode) to produce a fixed-length 300-dimensional document vector. That vector is passed through a two-layer MLP head (300→256 with ReLU and 0.3 dropout, then 256→4), yielding a 4-dimensional logit vector, and at inference the predicted label is the index of the largest logit. The unperturbed mechanism achieves $0.9705$ accuracy.
@@ Appendix A Automatic Efficient PAC Privatization
Recent work by Sridhar et al. [43] addresses these limitations through Efficient-PAC (Algorithm 3), which introduces two key improvements. First, they develop an anisotropic noise calibration scheme that avoids full covariance estimation by projecting mechanism outputs onto a unitary basis and estimating only per-direction variances. This leads to a more scalable and sample-efficient algorithm while maintaining rigorous mutual information guarantees. Second, they propose methods for reducing output instability through regularization and canonicalization techniques, enabling more consistent noise calibration and better overall utility. These refinements are particularly impactful in high-dimensional or structure-sensitive learning tasks, where the original PAC scheme may incur unnecessary noise due to variability not intrinsic to the learning objective.
Algorithm 3 Efficient-PAC [43]
@@
Efficient-PAC induces additional conservativeness relative to Auto-PAC. When Efficient-PAC enforces $\mathrm{MI}(X;\mathcal{M}(X)+B)\leq\beta$, the proof of Theorem 6 in [43] (Theorem 1) yields
@@ Appendix B Membership Inference Attack
We first recall the standard definition of membership inference attacks formalized to match PAC Privacy [46, 43].
Definition 10 (Membership Inference Attack [46, 43]).
@@
Consider the per–record “name–and–shame” mechanism $M$ that, independently for each index $i$, outputs $(i,x_{i})$ with probability $p$ and $\bot$ otherwise. Let $x,x\textquoteright$ be neighboring databases that differ only in record $i$ with $x_{i}\neq x\textquoteright_{i}$, and define the event $E={(i,x_{i})}$. Then $\Pr[M(x)\in E]=p,\ \Pr[M(x\textquoteright)\in E]=0.$ The $(\epsilon,\bar{\delta})$–DP inequality for $E$ reads $p\leq e^{\epsilon}\cdot 0+\bar{\delta}=\bar{\delta}$, hence any $(\epsilon,\bar{\delta})$ satisfied by $M$ must obey $\bar{\delta}\geq p$. In particular, with the standard regime $\bar{\delta}\ll 1/n$ (negligible failure probability), such a mechanism is not DP for any finite $\epsilon$; conversely, allowing $\bar{\delta}\geq p$ makes the guarantee vacuous on the pp–fraction of runs that reveal $(i,x_{i})$ exactly.
@@ C.2 Fair Comparison Under MIA
DP can be understood through the lens of membership inference success rates. Consider the membership inference scenario from Definition 10, where we have a dataset of size $n=\frac{N}{2}$ (i.e., each individual data record has a $50\%$ probability of being included in the selected subset $X$). If a mechanism $\mathcal{M}$ is $(\epsilon,\bar{\delta})$-DP, then by [28, 25], an adversary’s ability to successfully infer whether a specific individual record $i$ is included in the dataset (i.e., posterior success rate $p_{o}=1-\delta_{i}$) is fundamentally limited:
@@
In addition, there is a relationship between the posterior success rate $p_{o}$ and the mutual information [43] (derived from (1)):
@@ Appendix G Proof of Theorem 3
$\widehat{\mathtt{D}}_{\mathrm{LCE}}\equiv\Big{[}\widehat{\mathcal{J}}\big{(}\widehat{f}_{\mathrm{tr}};S_{P}^{\mathrm{val}},S_{Q}^{\mathrm{val}}\big{)}-\Gamma_{\hat{\delta}}\Big{]}_{+}$.
@@ Appendix H Proof of Theorem 4
$\widehat{\mathtt{D}}_{Z}\equiv\frac{1}{2\lambda_{\max}(\Sigma_{Z})}\,\mathrm{SW}_{2}^{2}\big{(}P_{\mathcal{M},B},\widetilde{Q}_{\mathcal{M}}\big{)}$,
@@
In addition, it is well known (see e.g., [3]) that the sliced Wasserstein distance provides a lower bound on the true 2-Wasserstein distance:
@@
Then, the Talagrand inequality [45, 36]implies
@@ Appendix I Proof of Corollary 2
By Theorem 1 of [46], a mechanism $\mathcal{M}$ satisfies $(\delta,\rho,\mathcal{D})$-PAC privacy where
@@
    The gradient of Gaussian entropy is $\nabla_{\Sigma_{B}}H(\mathcal{N}(0,\Sigma_{Z}))=\frac{1}{2}\Sigma_{Z}^{-1}$. By de Bruijn’s identity [44],
@@
For the stationarity condition w.r.t. each $\sigma_{w}^{2}$ we need the gradient of mutual information. By [37], we have
@@
these equalities can hold only for as long as the required mutual information reduction does not exceed $\beta_{\mathtt{lab}}$. Therefore, $\sigma_{v}^{2}=0$ for every $v\in S_{\mathtt{lab}}$. With those label-directions undisturbed, each class margin $e_{\ell}-e_{j}$ retains its sign, whence $\arg\max_{i}(Z_{i}+B^{*}_{i})=\hat{y}$.
@@
where $\mathtt{Gap}_{\mathtt{d}}(\beta)=\mathtt{D}_{\mathrm{KL}}(P_{\mathcal{M},B_{\mathrm{PAC}}}\|\widetilde{Q}_{\mathcal{M}})\geq 0$. When $\mathcal{M}(X)$ is non-Gaussian, $\mathtt{Gap}_{\mathtt{d}}(\beta)>0$ for all $\beta>0$. By de Bruijin’s idensity (e.g., [38]),
@@
We use the Likelihood-Ratio Attack (LIRA) described in [6] to perform the empirical membership inference attacks (MIAs) on the mechanisms privatized by SR-PAC, Auto-PAC, Efficient-PAC, and DP in Section 6.2, using Iris and Rice datasets. The empirical posterior success rate (PSR) is measured as the average accuracy of the MIA.

### v3

@@ 4.1 Foundation of Residual-PAC Privacy
Consider any $f$-divergence $\mathtt{D}_{f}$, Theorem 1 of [xiao2023pac] shows that the posterior advantage $\Delta_{f}^{\delta}$ is bounded by the minimum $f$-divergence between the joint distribution of $(X,\mathcal{M}(X))$, denoted by $P_{X,\mathcal{M}(X)}$, and the product of the marginal distribution $P_{X}$ and any auxiliary output distribution $P_{W}$ independent of $X$:
@@
Corollary 2 follows from Theorem 1 of [xiao2023pac] and establishes that when $\mathcal{H}(X)$ is finite, residual privacy $\mathtt{R}_{f}^{\delta}$ is lower bounded by $\mathcal{H}(X|\mathcal{M}(X))-\mathtt{V}$, where $\mathtt{V}$ is independent of both data distribution $\mathcal{D}$ and mechanism $\mathcal{M}$. Since $\mathtt{V}$ is constant, $\mathtt{R}_{f}^{\delta}-\mathtt{V}$ effectively provides a privacy quantification lower-bounded by conditional entropy $\mathcal{H}(X|\mathcal{M}(X))$. More generally, inequality (11) holds without requiring $\mathcal{H}(X)<\infty$, provided $\mathtt{D}_{f}(\mathcal{D}\|\mathcal{R})<\infty$.
@@ 5.2 Directional-Selectivity of SR-PAC
Let $Z\in\mathbb{R}^{d}$ be an output vector produced by a deterministic mechanism $\mathcal{M}(X)$; throughout we assume $\Sigma_{Z}\succ 0$ and finite differential entropy $\mathcal{H}(Z)$. For any application, let $S_{\mathrm{task}}\subseteq\mathbb{R}^{d}$ denote a practitioner‑chosen task‑critical sub‑space (the directions whose preservation matters most) and write $\Pi_{\mathrm{task}}$ for the orthogonal projector onto it.
Classification tasks. In what follows we illustrate the theory with multi‑class classification, where $Z$ is the logit vector, $\hat{y}=\arg\max_{i}Z_{i}$, and $S_{\mathtt{lab}}:=\operatorname{span}\{e_{\ell}-e_{j}:\,j\neq\ell\}$, where $\mathtt{lab}$ means "label". Let $\Pi_{\mathtt{lab}}$ be the projector onto $S_{\mathtt{lab}}$. The analysis for a general $S_{\mathrm{task}}$ is identical after replacing $\mathtt{lab}$ by $\mathrm{task}$.
@@
For every unit vector $w$, let $g(w)\equiv\frac{1}{2}\mathtt{mmse}(\langle Z,w\rangle)$, where $\mathtt{mmse}(\langle Z,w\rangle)\equiv\mathbb{E}\Bigl[\bigl\langle Z,w\bigr\rangle\;-\;\mathbb{E}\bigl[\langle Z,w\rangle|Y\bigr]\Bigr]^{2}$ is the minimum mean‐squared error of estimating the scalar random variable $\langle Z,w\rangle$ from the noisy observation $Y=Z+B$.
@@ 6 Experiments
We conduct two sets of experiments to evaluate our approach. First, we compare SR-PAC against Auto-PAC and Efficient-PAC (Appendix A) using CIFAR-10 [Krizhevsky09learningmultiple], CIFAR-100 [Krizhevsky09learningmultiple], MNIST [lecun2010mnist], and AG-News [zhang2015character] datasets, with results presented in Section 6.1. We use (R-)PAC to refer to the family of SR-PAC, Auto-PAC, and Efficient-PAC. Second, we extend this comparison to include DP by equalizing optimal posterior success rates of membership inference (Appendix C.2) across all four methods (SR-PAC, Auto-PAC, Efficient-PAC, and DP), making their privacy budgets comparable. For this comparison, we use Iris [fisher1988iris] and Rice [cinar2019classification] datasets, with results shown in Section 6.2. All experiments focus on output perturbation. Appendix S provides more details of the experiments.
@@
CIFAR-100 and Base Classifier. CIFAR-100 contains 50,000 training and 10,000 testing color images (each $32\times 32\times 3$), equally divided among 100 fine‐grained classes (500 training and 100 testing images per class). Each image is converted to a $3\times 32\times 32$ tensor and normalized per channel to mean 0.5 and standard deviation 0.5 before being fed into the network. The unperturbed classifier is a deep convolutional neural network with three convolutional “blocks.” Each block consists of two $3\times 3$ convolutions (with BatchNorm and ReLU after each), followed by a $2\times 2$ max‐pool, which sequentially maps inputs from $32\times 32$$16\times 16$$8\times 8$$4\times 4$, with channel widths increasing from 3 → 64 → 128 → 256. After flattening the resulting $256\times 4\times 4$ feature map into a 4096‐dimensional vector, a three‐layer MLP head (4096→512→256→100) with ReLU activations and 0.5 dropout between the first two fully connected layers produces a 100‐dimensional logit vector. During training, this network minimizes cross‐entropy loss over the CIFAR-100 classes; at inference, each normalized image is mapped to its 100-dimensional logits, and the predicted label is given by the $\arg\max$ of those logits. The unperturbed classifier achieves $0.5914\pm 0.0090$ accuracy.
@@
AG-News dataset and Base Classifier. AG-News comprises 120,000 training and 7,600 test articles equally divided among four classes (World, Sports, Business, Sci/Tech), i.e., 30,000 training and 1,900 test examples per class. Each example’s title and description are concatenated into one text string, then lowercased and split on whitespace (truncated or padded to 64 tokens). We build a 30,000‐word vocabulary from the training split and map each token to its index (with out‐of‐vocabulary tokens as 0). Those indices feed into an nn.EmbeddingBag layer (embedding size 300, mean‐pooling mode) to produce a fixed‐length 300‐dimensional document vector. That vector is passed through a two‐layer MLP head (300→256 with ReLU and 0.3 dropout, then 256→4), yielding a 4‐dimensional logit vector, and at inference the predicted label is the index of the largest logit. The unperturbed mechanism achieves $0.9705$ accuracy.
@@ Appendix A Automatic Efficient PAC Privatization
Recent work by Sridhar et al. [sridhar2024pac] addresses these limitations through Efficient-PAC (Algorithm 3), which introduces two key improvements. First, they develop an anisotropic noise calibration scheme that avoids full covariance estimation by projecting mechanism outputs onto a unitary basis and estimating only per-direction variances. This leads to a more scalable and sample-efficient algorithm while maintaining rigorous mutual information guarantees. Second, they propose methods for reducing output instability through regularization and canonicalization techniques, enabling more consistent noise calibration and better overall utility. These refinements are particularly impactful in high-dimensional or structure-sensitive learning tasks, where the original PAC scheme may incur unnecessary noise due to variability not intrinsic to the learning objective.
Algorithm 3 Efficient-PAC [sridhar2024pac]
@@
Efficient-PAC induces additional conservativeness relative to Auto-PAC. When Efficient-PAC enforces $\mathrm{MI}(X;\mathcal{M}(X)+B)\leq\beta$, the proof of Theorem 6 in [sridhar2024pac] (Theorem 1) yields
@@ Appendix B Membership Inference Attack
We first recall the standard definition of membership inference attacks formalized to match PAC Privacy [xiao2023pac, sridhar2024pac].
Definition 10 (Membership Inference Attack [xiao2023pac, sridhar2024pac]).
@@
Consider the per–record “name–and–shame” mechanism $M$ that, independently for each index $i$, outputs $(i,x_{i})$ with probability $p$ and $\bot$ otherwise. Let $x,x$’ be neighboring databases that differ only in record $i$ with $x_{i}\neq x’_{i}$, and define the event $E={(i,x_{i})}$. Then $\Pr[M(x)\in E]=p,\ \Pr[M(x’)\in E]=0.$ The $(\epsilon,\bar{\delta})$–DP inequality for $E$ reads $p\leq e^{\epsilon}\cdot 0+\bar{\delta}=\bar{\delta}$, hence any $(\epsilon,\bar{\delta})$ satisfied by $M$ must obey $\bar{\delta}\geq p$. In particular, with the standard regime $\bar{\delta}\ll 1/n$ (negligible failure probability), such a mechanism is not DP for any finite $\epsilon$; conversely, allowing $\bar{\delta}\geq p$ makes the guarantee vacuous on the pp–fraction of runs that reveal $(i,x_{i})$ exactly.
@@ C.2 Fair Comparison Under MIA
DP can be understood through the lens of membership inference success rates. Consider the membership inference scenario from Definition 10, where we have a dataset of size $n=\frac{N}{2}$ (i.e., each individual data record has a $50\%$ probability of being included in the selected subset $X$). If a mechanism $\mathcal{M}$ is $(\epsilon,\bar{\delta})$-DP, then by [kairouz2015composition, humphries2023investigating], an adversary’s ability to successfully infer whether a specific individual record $i$ is included in the dataset (i.e., posterior success rate $p_{o}=1-\delta_{i}$) is fundamentally limited:
@@
In addition, there is a relationship between the posterior success rate $p_{o}$ and the mutual information [sridhar2024pac] (derived from (1)):
@@ Appendix G Proof of Theorem 3
$\widehat{\mathtt{D}}_{\mathrm{LCE}}\equiv\Big[\widehat{\mathcal{J}}\big(\widehat{f}_{\mathrm{tr}};S_{P}^{\mathrm{val}},S_{Q}^{\mathrm{val}}\big)-\Gamma_{\hat{\delta}}\Big$$_{+}$.
@@ Appendix H Proof of Theorem 4
[…]
@@
In addition, it is well known (see e.g., [bonneel2015sliced]) that the sliced Wasserstein distance provides a lower bound on the true 2-Wasserstein distance:
@@
Then, the Talagrand inequality [talagrand1996transportation, otto2000generalization]implies
@@ Appendix I Proof of Corollary 2
By Theorem 1 of [xiao2023pac], a mechanism $\mathcal{M}$ satisfies $(\delta,\rho,\mathcal{D})$-PAC privacy where
@@
The gradient of Gaussian entropy is $\nabla_{\Sigma_{B}}H(\mathcal{N}(0,\Sigma_{Z}))=\frac{1}{2}\Sigma_{Z}^{-1}$. By de Bruijn’s identity [stam1959some],
@@
For the stationarity condition w.r.t. each $\sigma_{w}^{2}$ we need the gradient of mutual information. By [palomar2005gradient], we have
@@
these equalities can hold only for as long as the required mutual information reduction does not exceed $\beta_{\mathtt{lab}}$. Therefore, $\sigma_{v}^{2}=0$ for every $v\in S_{\mathtt{lab}}$. With those label‑directions undisturbed, each class margin $e_{\ell}-e_{j}$ retains its sign, whence $\arg\max_{i}(Z_{i}+B^{*}_{i})=\hat{y}$.
@@
where $\mathtt{Gap}_{\mathtt{d}}(\beta)=\mathtt{D}_{\mathrm{KL}}(P_{\mathcal{M},B_{\mathrm{PAC}}}\|\widetilde{Q}_{\mathcal{M}})\geq 0$. When $\mathcal{M}(X)$ is non-Gaussian, $\mathtt{Gap}_{\mathtt{d}}(\beta)>0$ for all $\beta>0$. By de Bruijin’s idensity (e.g., [park2012equivalence]),
@@
We use the Likelihood-Ratio Attack (LIRA) described in [carlini2022membership] to perform the empirical membership inference attacks (MIAs) on the mechanisms privatized by SR-PAC, Auto-PAC, Efficient-PAC, and DP in Section 6.2, using Iris and Rice datasets. The empirical posterior success rate (PSR) is measured as the average accuracy of the MIA.